OpenStudy (anonymous):

Solve exactly. log(base 2)(23-x)=2 x= ?

OpenStudy (anonymous):

y = log (base b)x is equivalent to b^y = x

OpenStudy (anonymous):

Think I got this one as the answer I came up with was 19.

OpenStudy (anonymous):

I am stuck on x^2=log(x+5) however

OpenStudy (anonymous):

and yes I used what you typed Denebal thank you